HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

DUAL-COMB SPECTROSCOPY IN THE TWO-MICRON RANGE USING A NOVEL DESIGN OF DISPERSION-CONTROLLED HIGHLY NONLINEAR

FIBRE

\underline{A. MALFONDET} , M. DEROH, G. MILLOT, ICB, Université de Bourgogne, CNRS UMR 6303, F-21078 Dijon, France,; S.-E. AHMEDOU, R. DAULIAT, P. ROY, R. JAMIER, Université de Limoges, CNRS, XLIM, UMR 7552, F-87000 Limoges, France; A. PARRIAUX, Laboratoire Temps-Fréquence, Institut de Physique, Université de Neuchâtel, Avenue de Bellevaux 51, 2000 Neuchâtel,

Switzerland; L. LABONTE, S. TANZILLI, Univ. Côte d’Azur, CNRS, Institut de Physique de Nice, F-06108 Nice Cedex 2, France; J.-C. DELAGNES, CELIA, Centre Laser Intenses et Applications, Université de Bordeaux-CNRS-CEA, UMR

5107, F-33405 Talence Cedex France

Over the past few decades, dual-comb spectroscopy (DCS) has garnered significant interest due to its ability to enable precise and real-time gas detection ^{1}. This technique finds applications in various fields such as breath analysis, air pollutant detection, agricultural gas flux measurement, or material characterization to cite a few. Moreover, the availability of mature components operating at 1.55 μm has greatly facilitated the development of spectrometers based on DCS. However, certain molecular species, like CO_{2}, exhibit a very low absorption in this spectral window, necessitating long absorption lengths to detect them effectively. Hence, there is a need to convert the frequency combs to a more suitable spectral window, preferably around 2 μm or beyond, to combine the component maturity at 1.55 μm with efficient absorption in the mid-infrared region.

In this study, we employ the phenomenon of degenerate four-wave mixing ^{2} in a specially designed dispersion-controlled highly nonlinear fiber to convert our frequency combs in the two-micron region ^{3}. This approach enables us to measure the rotational-vibrational absorptions of ^{12}C^{16}O_{2} and ^{14}N_{2}^{16}O and extract the selfbroadening coefficients associated with these molecules. A study is also performed on a mixture^{12}C^{16}O_{2}/^{14}N_{2}^{16}O, in order to measure the broadening of the CO_{2} peaks due to the presence of N_{2}O. Our results are in very good agreement with the HITRAN database, showing the performance of the proposed dual-comb spectrometer.

1G. Millot et al., "Frequency-agile dual-comb spectroscopy", Nature Photonics, 10, 27 (2016)

2S. Pitois et al.., "Experimental observation of a new modulation instability spectral window induced by fourth-order dispersion in a normally dispersive single-mode optical fiber", Optics Communications, 226,415-422 (2003)

3S.E. Ahmedou et al., "Design and fabrication of dispersion controlled highly nonlinear fibers for far-detuned four-wave mixing frequency conversion",Optics Express, 30, 6 (2023)

p-number: p220  Submitted on Sat, 17 Jun 2023 18:36:25 +0200